OpenStudy (bananas1234):

Multiply and simplify: (2x + 5)(3x + 4)

OpenStudy (anonymous):

do you know how to multiply and add? \((ax + c) (bx + d) = ax \times bx + ax \times d + bx \times c + c \times d \)

OpenStudy (bananas1234):

i am not very good at it

OpenStudy (anonymous):

http://www.ck12.org/arithmetic/

OpenStudy (anonymous):

http://www.ck12.org/algebra/

OpenStudy (texaschic101):

FOIL method (2x + 5)(3x - 4) F -- multiply first terms in each set....2x * 3x = ? O -- multiply outside terms in each set ... 2x * - 4 = ? I -- multiply inside terms in each set...5 * 3x = ? L -- multiply last terms in each set...5 * -4 = ? Now after you do the multiplying, what do you end up with ?

OpenStudy (bananas1234):

so 2x * 3x = 6x^2

OpenStudy (texaschic101):

oops...I wrote the problem wrong... you basically do the same a far as the multiplying...just use the right numbers (2x + 5)(3x + 4) now do FOIL

OpenStudy (texaschic101):

you got the 6x^2 right....thats the F....now do the O....outside terms 2x * 4 = ?

OpenStudy (bananas1234):

and 2x * 4 = 8x?

OpenStudy (texaschic101):

correct... now do the Inside terms of each set....5 * 3x = ?

OpenStudy (bananas1234):

15x

OpenStudy (texaschic101):

Now do the L....last terms in each set ...5 * 4 = ?

OpenStudy (bananas1234):

20

OpenStudy (texaschic101):

now put all the answers you got together.. 6x^2 + 8x + 15x + 20 now just combine like terms and what do you get ?

OpenStudy (bananas1234):

6x2 + 23x + 20

OpenStudy (texaschic101):

correct...thats it

OpenStudy (bananas1234):

Awesome! Thank you very much! :)

OpenStudy (texaschic101):

no problem :)